Three angles of a triangle ABC are in Arithmetic progression and two sides are in the ratio b : c = √3 : √2. Find angle A.

Grade:

1 Answers

Sourabh Singh IIT Patna
9 years ago

Hii




Use property of AP and then use the formulae of cosine from solution of triangles to get the answer in the required format.

[a^2 = b^2 + c^2 - 2bc\cos\alpha\,]





Best